A 10 m wide spacecraft moves through the interstellar space at a speed 3 × 10^7 m/s. A magnetic field B = 3 × 10^-10 T exists in the space in a direction perpendicular to the plane of motion. Treating the spacecraft as a conductor, calculate the emf induced across its width.

Grade:upto college level

1 Answers

Kevin Nash
9 years ago
Sol. ℓ = 10 m, v = 3 × 10^7 m/s, B = 3 × 10^–10 T Motional emf = Bvℓ = 3 × 10^–10 × 3 × 10^7 × 10 = 9 × 10^–3 = 0.09 V